Find each of the following squares.
(7p+6q)^2
Find each of the following squares.
(w minus(1)divided by (4))^2
Find each of the following products.
a(a+5)(a-5)

Answer by algebrapro18(249)

the answers will be in red

1)Find each of the following squares.
(7p+6q)^2

(7p+6q)^2=(7p+6q)(7p+6q)

now just use the foil method

f = 7p(7p)
o = 7p(6q)
i = 7p(6q)
l = 6q(6q)

(7p+6q)(7p+6q)
=7p(7p)+7p(6q)+7p(6q)+6q(6q)
=49p^2+42pq+42pq+36q^2
=9p^2-84pq+36q

2)Find each of the following squares.
(w minus(1)divided by (4))^2

(w minus(1)divided by (4))^2 = (w-1)^2/4^2

(w-1)^2 = (w-1)(w-1)

we use foil again on the top

f = w(w)
o = w(-1)
i = -1(w)
l = -1(-1)

(w-1)(w-1)
= w(w)+w(-1)+-1(w)+-1(-1)
=w^2-w-w+1
=w^2-2w+1

(w-1)^2/4^2
=(w^2-2w+1)/16

3)Find each of the following products.
a(a+5)(a-5)

use foil first

f = a(a)
o = a(-5)
i = 5(a)
l = 5(-5)

(a+5)(a-5)
= a(a)+a(-5)+5(a)+5(-5)
= a^2-5a+5a-25
=a^2-25
now distribute
a(a+5)(a-5)
=a(a^2-25)
=a^3-25a